x+2y=6
x-3y=-4
Help me solve thiss
!!

Answer by fractalier(2101)

From
x + 2y = 6
x - 3y = -4
we can subtract equations and eliminate the x's...we get
x + 2y = 6
-(x - 3y = -4)
and we have
5y = 10
y = 2
then
x = 2 also...